Evaluate the integral of (e^-x+e^x)² dx from 0 to 1


1
Expert's answer
2021-02-24T14:17:50-0500

Solution.

"\int\limits_0^1 (e^{-x}+e^x)^2dx=\int\limits_0^1(e^{-2x}+2e^{-x}e^x+e^{2x})dx=\newline
=(-\frac{1}{2}e^{-2x}+2x+\frac{1}{2}e^{2x})|_0^1=-\frac{1}{2}e^{-2}+2+\frac{1}{2}e^2+\frac{1}{2}
-\frac{1}{2}=\newline
=-\frac{1}{2e^2}+\frac{e^2}{2}+2."

Answer. "-\frac{1}{2e^2}+\frac{e^2}{2}+2."